On hot, sunny, summer days, Jane rents inner tubes by the river that runs through her town. Based on her past experience, she has assigned the following probability distribution to the number of tubes she will rent on a randomly selected day.

x 25 50 75 100 Total
P(x) .30 .36 .20 .14 1.00

Find the probability expressions: (Round your answers to 2 decimal places.

a. P(X = 25)
b. P(X ≤ 50)
c. P(X > 25)
d. P(X < 75)
Answer #1

a) the probability expression P(X = 25)  
   P(X = 25) = 0.30

b) the probability expression P(X ≤ 50)
P(X <= 50) = P(X=25) + P(50)
= 0.30 + 0.36 = 0.66
P(X ≤ 50) = 0.66
c) the probability expression P(X > 25)
   P(X > 25) = 1 - P(X<=25)
= 1 - 0.30

  P(X > 25) =0.70
  
d) the probability expression   P(X < 75)
P(X < 75) = 1 - P(X>=75)
= 1 - 0.20 - 0.14
P(X < 75) = 0.46